Write three whole numbers occurring just before 10001.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
10001 – 1 = 10000 10000 – 1 = 9999 9999 – 1 = 9998 Hence, three whole numbers just before 10001 are 10000, 9999 and 9998.